OpenStudy (anonymous):

Solve for x: (x+4)/2 = 7 x = 3 x = 5 x = 6 x = 10

jimthompson5910 (jim_thompson5910):

You're given a list of possible choices. So you can plug in each choice in for x and see if you get a true equation.

OpenStudy (chris911):

10 refer to https://answers.yahoo.com/question/index?qid=20111026104151AAHYU0L

OpenStudy (anonymous):

10

jimthompson5910 (jim_thompson5910):

For example, if x = 6, then... (x+4)/2 = 7 (6+4)/2 = 7 10/2 = 7 5 = 7 ... which is false, so x = 6 is not the solution

OpenStudy (anonymous):

Thank guys and @jim_thompson5910 that actually kinda makes sense haha

jimthompson5910 (jim_thompson5910):

I'm glad it does